# No Other Representation Component Is Needed: Diffusion Transformers Can Provide Representation Guidance by Themselves

### v1

@@ Abstract
Recent studies have demonstrated that learning a meaningful internal representation can both accelerate generative training and enhance generation quality of the diffusion transformers. However, existing approaches necessitate to either introduce an additional and complex representation training framework or rely on a large-scale, pre-trained representation foundation model to provide representation guidance during the original generative training process. In this study, we posit that the unique discriminative process inherent to diffusion transformers enables them to offer such guidance without requiring external representation components. We therefore propose Self-Representation Alignment (SRA), a simple yet straightforward method that obtain representation guidance through a self-distillation manner. Specifically, SRA aligns the output latent representation of the diffusion transformer in earlier layer with higher noise to that in later layer with lower noise to progressively enhance the overall representation learning during only generative training process. Experimental results indicate that applying SRA to DiTs and SiTs yields consistent performance improvements. Moreover, SRA not only significantly outperforms approaches relying on auxiliary, complex representation training frameworks but also achieves performance comparable to methods that heavily dependent on powerful external representation priors. Refer to caption Figure 1: Left: Methods like MaskDiT [77] and SD-DiT [79] use an additional representation task to regulate diffsuion transformer. Middle: Methods like REPA [74] leverage an additional representation foundation model as guidance. Right (our approach): We do not use any representation component but only seek to provide representation guidance by diffusion transformer itself.
@@ 1 Introduction
Recently, many works [77, 79, 74, 39] have explored leveraging representation components of vision transformers for diffusion transformer’s training, and have shown that learning a high-quality internal representation can not only speed-up the generative training progress, but also improve the generation quality. These works either utilize the training paradigm in representation learning (e.g., MAE’s [26], IBOT’s [78]) shown in Figure 1a or leverage a large-scale pre-trained representation foundation model (e.g., DINOv2 [56], CLIP [60]) shown in Figure 1b to give representation guidance for the diffusion transformer during the original generative training. However, the former method requires a complex training framework design, and the latter method relies on a powerful prior which is trained on massive amount of data with thousands of GPUs. Thus, we raise a significant but underexplored question: Can we obtain such representation guidance only during the generative training without any representation component involved?
@@
Figure 2: We empirically investigate the representations in diffusion transformers across different blocks and timesteps with the original SiT-XL/2 checkpoint trained for 7M iterations. Left: Using PCA [1] to visualize the latent features in SiT, we observe that the features lead a process from coarse to fine when increasing block layers and decreasing noise level. Right: The similar trend can also be seen in the linear probing results on ImageNet. Investigation of DiT is provided in Appendix A, which leads to the similar results as SiT.
@@
Our observations: Different from the representation model that takes a clean image as input then output semantically-rich feature, diffusion model often takes a noise latent as input and obtain cleaner one step-by-step. In other words, the generative mechanism by which the diffusion model operates can be generally considered as a coarse to fine process. Hence, we hypothesise that the representations in diffusion models also follow this trend. To testify this, we perform an empirical analysis with recent diffusion transformers [52, 57]. As shown in Figure 2a, we first find out that the latent features in the diffusion transformer are progressively refined, moving from coarse to fine, as block layers increase and noise level decreased. Next, akin to the results in prior studies [74, 70], we observe that the diffusion transformer already learns meaningful discriminative representations as shown in Figure 2b. Meanwhile, although the accuracy drop off after reach a peak at about layer 20 because the model needs to shift away to focus on generating images with high-frequency details, the quality of the representations basically transfer from bad to good by increasing block layers and decreasing noise level. These results indicate that the diffusion transformer gets a roughly from coarse-to-fine discriminative process when only generative training is performed.
@@
This distinctive trend motivates us to leverage the better representations in the diffusion transformer to guide the weaker ones in the original generative training, thereby enhancing the representation learning of the model without involving any additional representation component. However, this approach is not straightforward during training. As observed in previous studies [11, 25], directly using the output of the same model as a supervision signal may lead to shortcut learning [23], which can compromise performance or even cause training collapse.
@@
Our approach: To overcome abovementioned hurdles, we present Self-Representation Alignment (SRA), a simple but effective technique built on recent diffusion transformer architectures [52, 57]. As shown in Figure 1c, SRA dose not need any representation component, in essence, it aligns the output latent representations in earlier layer conditioned with higher noise to that in later layer conditioned with lower noise to achieve self-representation guidance. Meanwhile, in order to make the training process more stable, we obtain the target features from another model which shares the same architecture with the trainable model but updates weight by weighted moving average (EMA)***In the following part, we use term ‘student’ to represent the trainable model and ‘teacher’ to represent the EMA model in SRA for simplicity.. Furthermore, the student’s output latent feature is first passed through the projection layers to conduct a slight nonlinear transformation for better representation extraction, and then aligned with the target feature output by the teacher. In a nutshell, our SRA can offer a flexible way to integrate representation guidance without representation component needs and architect modification.
@@
Based on our analysis, we conduct comprehensive experiments to evaluate the effectiveness of SRA. After a series of component-wise analyses within the design space, we show that SRA brings significant performance improvements to both DiTs [57] and SiTs [52]. Moreover, our ablation study highlights the crucial role of internal representations in the success of SRA, which supports our central hypothesis: diffusion transformers can provide representation guidance by themselves.
@@
  • We analyze the latent representations in the diffusion transformers and assume that diffusion transformers have the potential to provide representation guidance by themselves.
  • We introduce SRA, a simple yet effective method that aligns the output latent representations of the diffusion transformers in earlier layer with higher noise to that in later layer with lower noise to achieve self-representation guidance.
@@ 2.1 Preliminary: Training Object of DiT and SiT
As our method is built upon the diffusion-based model (DiT) and flow-based model (SiT), to facilitate a more seamless introduction of SRA in the subsequent part, we now present a brief overview of the training object of these two types of models. We omit the class-condition here for simplicity because it can be easily added to all analogs without causing any changes to the mathematical equations. We leave more detailed mathematical descriptions of these types of models in Appendix B.
@@
Diffusion-based models learn to transform Gaussian noise into data samples through a step-by-step denoising process. Given a pre-defined forward process that gradually adds noise, these models learn the reverse process to recover the original data.
@@
For data point ${\mathbf{x}}_{0}$ from distribution $p({\mathbf{x}})$, the forward noising process follows: $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{1-\beta_{t}}{\mathbf{x}}_{0},\beta_{t}^{2}\mathbf{I})$. The model learns to reverse this process using a neural network $\bm{\epsilon}_{\bm{\theta}}({\mathbf{x}}_{t},t)$ that predicts the noise added at each step. The network is trained using a simple mean squared error objective that measures how well it can predict the noise:
$\mathcal{L}_{\text{simple}}=\mathbb{E}{{\mathbf{x}}_{\ast},\bm{\epsilon},t}\Big{[}||\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}({\mathbf{x}}_{t},t)||_{2}^{2}\Big{]}.$ (1)
@@
Different from diffusion-based model, flow-based model aims to learn a velocity field ${\mathbf{v}}_{\theta}({\mathbf{x}}_{t},t)$ that governs a probability flow ordinary differential equation (PF ODE). This PF ODE allows the model to sample data by starting from Gaussian noise and flowing towards the data distribution. This process is defined as:
${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{\ast}+\sigma_{t}{\mathbf{\epsilon}},\quad\alpha_{0}=\sigma_{T}=1,\alpha_{T}=\sigma_{0}=0,$ (2)
@@
where ${\mathbf{x}}_{\ast}\sim p({\mathbf{x}})$ is the data, ${\mathbf{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is Gaussian noise, and $\alpha_{t}$ and $\sigma_{t}$ are decreasing and increasing functions of $t\in[0,T]$, respectively. The PF ODE is given by:
$\dot{{\mathbf{x}}}_{t}={\mathbf{v}}({\mathbf{x}}_{t},t),$ (3)
where the distribution of this ODE at time $t$ matches the marginal distribution $p_{t}({\mathbf{x}})$.
The velocity field can be expressed as:
${\mathbf{v}}({\mathbf{x}},t)=\mathbb{E}[\dot{{\mathbf{x}}}_{t}|{\mathbf{x}}_{t}={\mathbf{x}}]=\dot{\alpha}_{t}\mathbb{E}[{\mathbf{x}}_{\ast}|{\mathbf{x}}_{t}={\mathbf{x}}]+\dot{\sigma}_{t}\mathbb{E}[{\mathbf{\epsilon}}|{\mathbf{x}}_{t}={\mathbf{x}}].$ (4)
@@
To learn the velocity field, the model ${\mathbf{v}}_{\theta}({\mathbf{x}}_{t},t)$ is trained to minimize the following loss function:
@@
$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{{\mathbf{x}}_{\ast},\bm{\epsilon},t}\big{[}||{\mathbf{v}}_{\theta}({\mathbf{x}}_{t},t)-\dot{\alpha}_{t}{\mathbf{x}}_{\ast}-\dot{\sigma}_{t}{\mathbf{\epsilon}}||^{2}\big{]}.$ (5)
@@
For the sake of simplicity, in the following part, we use generative loss ($\mathcal{L}_{\text{gen}}$) to uniformly represent the two generative training objects of DiT and SiT.
@@ 2.2 Self-Representation Alignment
Figure 3: Overall framework of SRA. Our SRA explicitly aligns the diffusion transformer’s representations in earlier layer with higher noise with the representations in later layer with lower noise to conduct self-representation alignment.
@@
Previous studies have demonstrated that learning good internal representation can both speed up diffusion transformer’s training and improve the quality of its generated samples. In SRA, our insight is aligning student’s latent feature in earlier layer conditioned on higher noise with that in later layer conditioned on lower noise of teacher to conduct self-representation guidance without requiring any dedicated representation components. As depicted in Figure 3, the goal of this simple training framework is let the diffusion transformer not only predict noise-invariant, but also align with better visual representations from itself. This thereby provides a simple and resource-friendly way to enhance the representation learning in the diffusion transformer during generative training without the need to design complex representation regulation or introduce additional representation foundation models.
@@
Formally, let $f$ be the trainable student model and $f_{\ast}$ be the teacher model. Considering input noise latent, timestep, and condition to be ${\mathbf{x}}_{t}$, $t$, and $c$. Then, we can obtain the student encoder latent output ${\mathbf{y}}=f^{m}({\mathbf{x}}_{t},t,c)\in\mathbb{R}^{B\times N\times D}$, where $B,N,D>0$ are the batchsize, number of patches and the embedding dimension for $f$, and $m$ denotes the output from the $m^{th}$ layer in $f$. Similarly, the output of the teacher can be expressed as ${\mathbf{y}}_{\ast}=f_{\ast}^{n}({\mathbf{x}}_{t-k},t-k,c)\in\mathbb{R}^{B\times N\times D}$. In our SRA, we set $m\leq n$, $k\geq 0$, and $0\leq(t-k)<t_{max}$or SiTs, $t_{max}=1$ and for DiTs, $t_{max}=1000$. In practice, we truncate ($t-k$) to 0 if it is less than 0.. Hence, we can conduct self-representation alignment using the teacher’s output ${\mathbf{y}}_{\ast}$ and the student’s output variant $h_{\phi}({\mathbf{y}})\in\mathbb{R}^{B\times N\times D}$, where $h_{\phi}({\mathbf{y}})$ is a projection of the student encoder output ${\mathbf{y}}$ that through a lightweight trainable MLPs head $h_{\phi}$. Notably, this projection head can be discarded optionally after training, which enables SRA to provide guidance without altering any architecture in diffusion transformers.
@@
In particular, SRA attains self-alignment by minimizing the patch-wise distance between the teacher’s output (${\mathbf{y}}_{\ast}$) and the student’s output variant ($h_{\phi}({\mathbf{y}})$):
@@
$\mathcal{L}_{\text{sa}}(\theta_{s},\phi)=\mathbb{E}_{{\mathbf{x}}_{t},t,c}\Big{[}\frac{1}{N}\sum_{i=1}^{N}\mathrm{dist}({\mathbf{y}}_{\ast}^{[i]},h_{\phi}({\mathbf{y}}^{[i]}))\Big{]},$ (6)
@@
where $[i]$ is a patch index, $\mathrm{dist}(\cdot,\cdot)$ is a pre-defined distance calculation function, and $\theta_{s}$, $\phi$ is the parameters of student diffusion transformer and the projection head.
Finally, we add this objective in the original diffusion-based objective described in Section 2.1 and Appendix B for joint learning:
@@
where $\lambda>0$ is a hyperparameter that controls the trade-off between generation object and self-representation alignment object.

### v3

@@ Abstract
Recent studies have demonstrated that learning a meaningful internal representation can both accelerate generative training and enhance the generation quality of diffusion transformers. However, existing approaches necessitate to either introduce an external and complex representation training framework or rely on a large-scale, pre-trained representation foundation model to provide representation guidance during the original generative training process. In this study, we posit that the unique discriminative process inherent to diffusion transformers enables them to offer such guidance without requiring external representation components. We therefore propose Self-Representation Alignment (SRA), a simple yet straightforward method that obtains representation guidance through a self-distillation manner. Specifically, SRA aligns the output latent representation of the diffusion transformer in the earlier layer with higher noise to that in the later layer with lower noise to progressively enhance the overall representation learning during only the generative training process. Experimental results indicate that applying SRA to DiTs and SiTs yields consistent performance improvements. Moreover, SRA not only significantly outperforms approaches relying on auxiliary, complex representation training frameworks but also achieves performance comparable to methods that are heavily dependent on powerful external representation priors. Refer to caption Figure 1: Left: Methods like MaskDiT [84] and SD-DiT [86] use an external representation task to guide diffusion transformer. Middle: Methods like REPA [81] leverage an external representation foundation model as guidance. Right (our approach): We do not use any external representation component but still obtain such guidance through proposed self-representation alignment technique.
@@ 1 Introduction
Recently, many works [84, 86, 81, 42] have explored leveraging representation components of vision transformers for diffusion transformer’s training and have shown that learning a high-quality internal representation can not only speed up the generative training progress but also improve the generation quality. These works either utilize the training paradigm in representation learning (e.g., MAE’s [29], IBOT’s [85]) shown in Figure 1a or leverage a large-scale pre-trained representation foundation model (e.g., DINOv2 [61], CLIP [65]) shown in Figure 1b to give representation guidance for the diffusion transformer during the original generative training. However, the former method requires a complex training framework design, and the latter method relies on a powerful prior, which is trained on a massive amount of data with hundreds of GPUs. Thereout, an intriguing yet underexplored problem has come to light: Can we obtain such representation guidance only during the generative training without external representation components?
@@
Figure 2: We empirically investigate the representations in diffusion transformers across different blocks and timesteps with the original SiT-XL/2 checkpoint trained for 7M iterations. Left: Using PCA [1] to visualize the latent features in SiT, we observe that the features lead a process from coarse to fine when increasing block layers and decreasing noise level. Right: A similar trend can also be seen in the linear probing results on ImageNet. Investigation of DiT is provided in Appendix A, which leads to the similar results as SiT.
@@
Our observations: Different from the representation model that takes a clean image as input and then outputs semantically-rich feature, diffusion model often takes a noise latent as input and obtains cleaner one step-by-step. In other words, the generative mechanism by which the diffusion model operates can be generally considered as a coarse to fine process. Inspired by this behavior, we hypothesise that the representations in it also follow such a trend. To testify this, we perform an empirical analysis with recent diffusion transformers [56, 62]. As shown in Figure 2a, we first find out that the latent features in the diffusion transformer are progressively refined, moving from coarse to fine, as block layers increase and noise level decreases. Next, akin to the results in prior studies [81, 77], we observe that the diffusion transformer already learns meaningful discriminative representations as shown in Figure 2b. Meanwhile, although the accuracy drops off after reaching a peak at about layer 20 because the model needs to shift away to focus on generating images with high-frequency details, the quality of the representations basically transfers from bad to good by increasing block layers and decreasing noise level. These indicate that the diffusion transformer gets roughly from a coarse-to-fine discriminative process when only generative training is performed.
@@
This distinctive trend motivates us to align the weaker representations in the diffusion transformer to the better ones in the original generative training, thereby enhancing the representation learning of the model without involving any external representation component. However, this approach is not straightforward during training. As observed in previous studies [12, 28], directly using the output of the same model as a supervision signal may lead to shortcut learning [26], which can compromise performance or even cause training collapse.
@@
Our approach: To overcome the abovementioned hurdles, we present Self-Representation Alignment (SRA), a simple but effective technique built on recent diffusion transformer architectures [56, 62]. As shown in Figure 1c, SRA does not need any representation component; in essence, it aligns the output latent representation in earlier layer conditioned with higher noise to that in later layer conditioned with lower noise to achieve self-representation alignment. Meanwhile, in order to make the training process more stable, we obtain the target features from another model that shares the same architecture with the trainable model but updates weight by weighted moving average (EMA)***In the following part, we use term ‘student’ to represent the trainable model and ‘teacher’ to represent the EMA model in SRA for simplicity.. Furthermore, the student’s output latent feature is first passed through the projection layers to conduct a slight nonlinear transformation for better representation extraction and then aligned with the target feature output by the teacher. In a nutshell, our SRA can offer a flexible way to integrate representation guidance without external component needs and architect modification.
@@
Finally, we conduct comprehensive experiments to evaluate the effect of SRA. After a series of component-wise analyses within the design space, we show that SRA brings significant performance improvements to both DiTs [62] and SiTs [56]. Moreover, our ablation study highlights the crucial role of internal representations in the success of SRA, which supports our central hypothesis: diffusion transformers can achieve representation alignment without external components.
@@
  • We analyze the latent representation in the diffusion transformers and assume that the unique coarse to fine discriminative process makes it possible to achieve representation alignment without external components.
  • We introduce SRA, a simple yet effective method that aligns the output latent representation of the diffusion transformers in the earlier layer with higher noise to that in the later layer with lower noise to achieve self-representation alignment.
@@ 2.1 Preliminary: Training Object of DiT and SiT
As our method is built upon the denoise-based model (DiT) and flow-based model (SiT), to facilitate a more seamless introduction of SRA in the subsequent part, we now present a brief overview of the training object of these two types of models. We omit the class-condition here for simplicity because it can be easily added to all analogs without causing any changes to the mathematical equations. We leave more detailed mathematical descriptions of these types of models in Appendix B.
@@
Denoise-based models learn to transform Gaussian noise into data samples through a step-by-step denoising process. Given a pre-defined forward process that gradually adds noise, these models learn the reverse process to recover the original data.
@@
For data point ${\mathbf{x}}_{0}$ from distribution ${\mathbf{x}}_{0}\sim p({\mathbf{x}})$, the forward process follows: $q({\mathbf{x}}_{t}|{\mathbf{x}}_{t-1})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{1-\beta_{t}}{\mathbf{x}}_{0},\beta_{t}\mathbf{I})$. The model learns to reverse this process using a neural network $\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)$ that predicts the noise added at each step. The network is trained using a simple mean squared error objective that measures how well it can predict the noise:
$\mathcal{L}_{\text{simple}}=\mathbb{E}{{\mathbf{x}}_{t},\bm{\epsilon},t}\Big{[}||\bm{\epsilon}-\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)||_{2}^{2}\Big{]}.$ (1)
@@
Different from the denoise-based model, the flow-based model aims to learn a velocity field ${\mathbf{v}}_{\zeta}({\mathbf{x}}_{t},t)$ that governs a probability flow ordinary differential equation (PF ODE). This PF ODE allows the model to sample data by flowing toward the data distribution. This forward process of PF ODE is defined as:
${\mathbf{x}}_{t}=\alpha_{t}{\mathbf{x}}_{0}+\sigma_{t}{\mathbf{\epsilon}},\quad\alpha_{0}=\sigma_{T}=1,\alpha_{T}=\sigma_{0}=0,$ (2)
@@
where ${\mathbf{x}}_{0}\sim p({\mathbf{x}})$ is the data, ${\mathbf{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is Gaussian noise, and $\alpha_{t}$ and $\sigma_{t}$ are monotone decreasing and increasing functions of $t\in[0,T]$, respectively. The PF ODE is given by:
$\dot{{\mathbf{x}}}_{t}={\mathbf{v}}_{\bm{\zeta}}({\mathbf{x}}_{t},t),$ (3)
@@
Where the marginal distribution of this ODE at time $p_{t}({\mathbf{x}})$ matches the marginal distribution of the forward process. To learn the velocity field, the model ${\mathbf{v}}_{\zeta}({\mathbf{x}}_{t},t)$ is trained to minimize the following loss function:
@@
$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{{\mathbf{x}}_{0},\bm{\epsilon},t}\big{[}||{\mathbf{v}}_{\bm{\zeta}}({\mathbf{x}}_{t},t)-\dot{\alpha}_{t}{\mathbf{x}}_{0}-\dot{\sigma}_{t}\bm{{\mathbf{\epsilon}}}||^{2}\big{]}.$ (4)
@@
For the sake of simplicity, in the following part, we use generative loss ($\mathcal{L}_{\text{gen}}$) to uniformly represent the two generative training objectives of denoise-based and flow-based methods.
@@ 2.2 Self-Representation Alignment
Figure 3: Overall framework of SRA. SRA explicitly aligns the diffusion transformer’s representation in the earlier layer with higher noise to that in the later layer with lower noise to achieve self-representation alignment.
@@
Previous studies have demonstrated that learning good internal representation can both speed up the diffusion transformer’s training and improve the quality of its generated samples. In SRA, our insight is aligning the student’s latent feature in the earlier layer conditioned on higher noise with that in the later layer conditioned on the lower noise of the teacher to conduct self-representation alignment without requiring any external representation components. As depicted in Figure 3, the goal of our simple training framework is to let the diffusion transformer not only predict noise-invariant but also align with better visual representations from itself. This operation thereby provides a simple way to enhance the representation learning in the diffusion transformer during generative training without the need to design complex representation regulation or introduce external representation foundation models.
@@
Formally, let $f$ be the trainable student model and $f_{\ast}$ be the teacher model. Considering input noise latent, timestep, and condition to be ${\mathbf{x}}_{t}$, $t$, and $c$. Then, we can obtain the student encoder latent output ${\mathbf{y}}=f^{m}({\mathbf{x}}_{t},t,c)\in\mathbb{R}^{B\times N\times D}$, where $B,N,D>0$ are the batchsize, number of patches and the embedding dimension for $f$, and $m$ denote the output from the $m^{th}$ layer in $f$. Similarly, the output of the teacher can be expressed as ${\mathbf{y}}_{\ast}=f_{\ast}^{n}({\mathbf{x}}_{t-k},t-k,c)\in\mathbb{R}^{B\times N\times D}$. In our SRA, we set $m\leq n$, $k\geq 0$, and $0\leq(t-k)<t_{max}$or SiTs, $t_{max}=1$ and for DiTs, $t_{max}=1000$. In practice, we truncate ($t-k$) to 0 if it is less than 0.. Hence, we can conduct self-representation alignment using the teacher’s output ${\mathbf{y}}_{\ast}$ and the student’s output transform $j_{\psi}({\mathbf{y}})\in\mathbb{R}^{B\times N\times D}$, where $j_{\psi}({\mathbf{y}})$ is a projection of the student encoder output ${\mathbf{y}}$ that through a lightweight trainable MLPs head $j_{\psi}$. Notably, this projection head can be discarded optionally after training, which enables SRA to provide guidance without altering any architecture in diffusion transformers.
@@
In particular, SRA attains self-alignment by minimizing the patch-wise distance between the teacher’s output (${\mathbf{y}}_{\ast}$) and the student’s output variant ($j_{\psi}({\mathbf{y}})$):
@@
$\mathcal{L}_{\text{sa}}(\zeta_{s},\psi)=\mathbb{E}_{\bm{{\mathbf{x}}}_{t},t,c}\Big{[}\frac{1}{N}\sum_{i=1}^{N}\mathrm{dist}({\mathbf{y}}_{\ast}^{[i]},j_{\psi}({\mathbf{y}}^{[i]}))\Big{]},$ (5)
@@
where $[i]$ is a patch index, $\mathrm{dist}(\cdot,\cdot)$ is a pre-defined distance calculation function, and $\zeta_{s}$, $\psi$ is the parameters of student diffusion transformer and the projection head. This objective is similar to that mentioned in REPA [81], except that we’re aiming to align with the diffusion transformer itself.
Finally, we add the abovementioned two objectives for joint learning:
@@
where $\lambda>0$ is a hyperparameter that controls the trade-off between the generation object and the self-representation alignment object.